# Hai-Long Sun1,2,3   Da-Wei Zhou1,2  Yang Li3  Shiyin Lu3  Chao Yi1,2 Qing-Guo Chen3  Zhao Xu3  Weihua Luo3  Kaifu Zhang3  De-Chuan Zhan1,2  Han-Jia Ye1,2

### v1

@@ Abstract
The rapid development of Multimodal Large Language Models (MLLMs) like GPT-4V has marked a significant step towards artificial general intelligence. Existing methods mainly focus on aligning vision encoders with LLMs through supervised fine-tuning (SFT) to endow LLMs with multimodal abilities, making MLLMs’ inherent ability to react to multiple languages progressively deteriorate as the training process evolves. We empirically find that the imbalanced SFT datasets, primarily composed of English-centric image-text pairs, lead to significantly reduced performance in non-English languages. This is due to the failure of aligning the vision encoder and LLM with multilingual tokens during the SFT process. In this paper, we introduce Parrot, a novel method that utilizes textual guidance to drive visual token alignment at the language level. Parrot makes the visual tokens condition on diverse language inputs and uses Mixture-of-Experts (MoE) to promote the alignment of multilingual tokens. Specifically, to enhance non-English visual tokens alignment, we compute the cross-attention using the initial visual features and textual embeddings, the result of which is then fed into the MoE router to select the most relevant experts. The selected experts subsequently convert the initial visual tokens into language-specific visual tokens. Moreover, considering the current lack of benchmarks for evaluating multilingual capabilities within the field, we collect and make available a Massive Multilingual Multimodal Benchmark which includes 6 languages, 15 categories, and 12,000 questions, named as MMMB. Our method not only demonstrates state-of-the-art performance on multilingual MMBench and MMMB, but also excels across a broad range of multimodal tasks. Both the source code and the training dataset of Parrot will be made publicly available.
@@ 1 Introduction
The rapid development of Large Language Models (LLMs), such as GPT-4 [52; 8; 45; 46; 53; 59], has gained significant attention. However, LLMs are limited to processing a single textual modality. The expansion into visual modalities has endowed LLMs with multimodal capabilities [64; 2; 63; 74; 17; 27; 32; 65; 71; 7], thereby accelerating the development of Multimodal Large Language Models (MLLMs) and further bringing us closer to the realization of Artificial General Intelligence (AGI).
@@ 2 Related Work
Multimodal Large Language Models. The domain of MLLMs has witnessed significant advances, particularly in the enhancement of visual and language processing. Current MLLM is usually a combination of visual encoders [51; 57; 21; 69; 48; 67], LLMs, and fusion modules. Innovations like Flamingo [2] and OpenFlamingo [4] have advanced visual representation by integrating a Perceiver Resampler with vision encoders. BLIP-2 [31] and InstructBLIP [17] employ Q-Former to connect the frozen LLM and vision encoder. InternVL [14] trains huge ViT and QFormer to integrate visual modalities through a multi-stage training method. MiniGPT4 [74] leverages both a Q-Former and a linear projector to bridge the gap between the vision module and LLM. Furthermore, LLaVA [38] adopts a simple MLP projector to promote the alignment between the LLM and vision encoder. mPLUG-Owl [64] introduces an approach that begins to finetune the vision encoder and align visual features, followed by tuning the LLM using LoRA [23]. Qwen-VL [6] improves visual module resolution to 448, aiming to refine the model’s visual processing capabilities. Fuyu-8B [7] directly projects image patches before integration with LLM. MM1 [43] has conducted ablative studies on connector design choices, revealing that the modality adapter type is less critical than the number of visual tokens and the resolution. MiniGemini [34] utilizes high-resolution visual tokens and high-quality data to narrow the performance gap with GPT-4 and Gemini. With the rapid advancements in open-source models, proprietary models such as GPT-4V/4o [46; 47], Gemini [58; 54], Qwen-VL-Plus/MAX [6], and Claude3 [3] have achieved outstanding results in evaluations and practical applications. In this work, owing to the simplicity of the LLaVA architecture, we adopt a framework similar to LLaVA to design our model.
@@
Multilingual Multimodal Models. Recent years have witnessed rapid progress in the expansion of multimodal models to include a wider variety of languages. M3[44] leverages English as a pivot and alternates between English-only vision-language pre-training and multilingual masked language modeling. In contrast, UC2 [73] translates English captions into various languages and uses images as the anchor. mCLIP [10] enhances the CLIP model by aligning it with a multilingual text encoder through knowledge distillation. Thanks to the expansion of the overall capabilities of large language models [1; 5; 26; 66], their multilingual capacities have significantly improved. Integrating multilingual LLMs with visual abilities has increasingly become a research focus. In the domain of LLMs, PaLI [13] develops a 17B multilingual language-image model that spans over 100 languages. Ying-VLM [33] discovers that instruction tuning in English can extend its applicability to other languages. Ziya-Visual [40] illustrates the translation of English image-text datasets into Chinese, using in-context learning for instruction-response generation. VisCPM [24] introduces a training paradigm that fine-tunes the MLLM in a quasi-zero-shot manner based on a strong multilingual large language model. Despite these advancements, they are primarily confined to two languages or rely on the massive translated corpus. On the other hand, there is no suitable multilingual benchmark for MLLMs to evaluate the performance of multiple languages. There are also some multilingual research studies in other domains, such as multilingual machine translation [72; 49; 50; 68].
@@ 3.1 Limitations of Existing Benchmarks
Figure 2: Some bad cases for multilingual benchmarkperceive. Left: code reasoning is strongly related to English. Middle: logical reasoning is too challenging. Right: lack relevance between image and text.
@@ 5.2 Main Results
LLaVA-1.5 [36] Vicuna-7B 67.07 58.83 59.76 43.50 46.43 59.06 65.37 58.33 59.02 36.16 43.90 56.95
LLaVA-1.5 [36] Vicuna-13B 69.76 62.86 60.76 45.49 54.44 62.69 68.98 63.23 62.97 46.56 53.17 61.59
LLaVA-NeXT [37] Vicuna-7B 70.87 61.57 61.81 42.74 46.95 63.85 67.95 60.56 60.39 38.40 45.36 59.62
LLaVA-NeXT [37] Vicuna-13B 74.44 67.19 63.21 45.36 53.09 68.24 70.87 64.51 64.08 45.36 52.92 61.85

### v2

@@ Abstract
The rapid development of Multimodal Large Language Models (MLLMs) like GPT-4V has marked a significant step towards artificial general intelligence. Existing methods mainly focus on aligning vision encoders with LLMs through supervised fine-tuning (SFT) to endow LLMs with multimodal abilities, making MLLMs’ inherent ability to react to multiple languages progressively deteriorate as the training process evolves. We empirically find that the imbalanced SFT datasets, primarily composed of English-centric image-text pairs, lead to significantly reduced performance in non-English languages. This is due to the failure of aligning the vision encoder and LLM with multilingual tokens during the SFT process. In this paper, we introduce Parrot, a novel method that utilizes textual guidance to drive visual token alignment at the language level. Parrot makes the visual tokens condition on diverse language inputs and uses Mixture-of-Experts (MoE) to promote the alignment of multilingual tokens. Specifically, to enhance non-English visual tokens alignment, we compute the cross-attention using the initial visual features and textual embeddings, the result of which is then fed into the MoE router to select the most relevant experts. The selected experts subsequently convert the initial visual tokens into language-specific visual tokens. Moreover, considering the current lack of benchmarks for evaluating multilingual capabilities within the field, we collect and make available a Massive Multilingual Multimodal Benchmark which includes 6 languages, 15 categories, and 12,000 questions, named as MMMB. Our method not only demonstrates state-of-the-art performance on multilingual MMBench and MMMB, but also excels across a broad range of multimodal tasks. Code is available at: https://github.com/AIDC-AI/Parrot.
@@ 1 Introduction
The rapid development of Large Language Models (LLMs), such as GPT-4 [52; 8; 45; 46; 53; 59; 66], has gained significant attention. However, LLMs are limited to processing a single textual modality. The expansion into visual modalities has endowed LLMs with multimodal capabilities [64; 2; 63; 75; 17; 27; 32; 65; 72; 7], thereby accelerating the development of Multimodal Large Language Models (MLLMs) and further bringing us closer to the realization of Artificial General Intelligence (AGI).
@@ 2 Related Work
Multimodal Large Language Models. The domain of MLLMs has witnessed significant advances, particularly in the enhancement of visual and language processing. Current MLLM is usually a combination of visual encoders [51; 57; 21; 70; 48; 68], LLMs, and fusion modules. Innovations like Flamingo [2] and OpenFlamingo [4] have advanced visual representation by integrating a Perceiver Resampler with vision encoders. BLIP-2 [31] and InstructBLIP [17] employ Q-Former to connect the frozen LLM and vision encoder. InternVL [14] trains huge ViT and QFormer to integrate visual modalities through a multi-stage training method. MiniGPT4 [75] leverages both a Q-Former and a linear projector to bridge the gap between the vision module and LLM. Furthermore, LLaVA [38] adopts a simple MLP projector to promote the alignment between the LLM and vision encoder. mPLUG-Owl [64] introduces an approach that begins to finetune the vision encoder and align visual features, followed by tuning the LLM using LoRA [23]. Qwen-VL [6] improves visual module resolution to 448, aiming to refine the model’s visual processing capabilities. Fuyu-8B [7] directly projects image patches before integration with LLM. MM1 [43] has conducted ablative studies on connector design choices, revealing that the modality adapter type is less critical than the number of visual tokens and the resolution. MiniGemini [34] utilizes high-resolution visual tokens and high-quality data to narrow the performance gap with GPT-4 and Gemini. With the rapid advancements in open-source models, proprietary models such as GPT-4V/4o [46; 47], Gemini [58; 54], Qwen-VL-Plus/MAX [6], and Claude3 [3] have achieved outstanding results in evaluations and practical applications. In this work, owing to the simplicity of the LLaVA architecture, we adopt a framework similar to LLaVA to design our model.
@@
Multilingual Multimodal Models. Recent years have witnessed rapid progress in the expansion of multimodal models to include a wider variety of languages. M3[44] leverages English as a pivot and alternates between English-only vision-language pre-training and multilingual masked language modeling. In contrast, UC2 [74] translates English captions into various languages and uses images as the anchor. mCLIP [10] enhances the CLIP model by aligning it with a multilingual text encoder through knowledge distillation. Thanks to the expansion of the overall capabilities of large language models [1; 5; 26; 67], their multilingual capacities have significantly improved. Integrating multilingual LLMs with visual abilities has increasingly become a research focus. In the domain of LLMs, PaLI [13] develops a 17B multilingual language-image model that spans over 100 languages. Ying-VLM [33] discovers that instruction tuning in English can extend its applicability to other languages. Ziya-Visual [40] illustrates the translation of English image-text datasets into Chinese, using in-context learning for instruction-response generation. VisCPM [24] introduces a training paradigm that fine-tunes the MLLM in a quasi-zero-shot manner based on a strong multilingual large language model. Despite these advancements, they are primarily confined to two languages or rely on the massive translated corpus. On the other hand, there is no suitable multilingual benchmark for MLLMs to evaluate the performance of multiple languages. There are also some multilingual research studies in other domains, such as multilingual machine translation [73; 49; 50; 69].
@@ 3.1 Limitations of Existing Benchmarks
Figure 2: Some bad cases for multilingual benchmark. Left: code reasoning is strongly related to English. Middle: logical reasoning is too challenging. Right: lack relevance between image and text.
@@ 5.2 Main Results
LLaVA-1.5 [36] Vicuna-v1.5-7B 67.07 58.83 59.76 43.50 46.43 59.06 65.37 58.33 59.02 36.16 43.90 56.95
LLaVA-1.5 [36] Vicuna-v1.5-13B 69.76 62.86 60.76 45.49 54.44 62.69 68.98 63.23 62.97 46.56 53.17 61.59
LLaVA-NeXT [37] Vicuna-v1.5-7B 70.87 61.57 61.81 42.74 46.95 63.85 67.95 60.56 60.39 38.40 45.36 59.62
LLaVA-NeXT [37] Vicuna-v1.5-13B 74.44 67.19 63.21 45.36 53.09 68.24 70.87 64.51 64.08 45.36 52.92 61.85